# Multimodal large language models and physics visual tasks: comparative analysis of performance and costs

### v1

@@ Abstract
Multimodal large language models (MLLMs) capable of processing both text and visual inputs are increasingly being explored for uses in physics education, such as tutoring, formative assessment, and grading. This study evaluates a range of publicly available MLLMs on a set of standardized, image-based physics research-based conceptual assessments (concept inventories). We benchmark 15 models from three major providers (Anthropic, Google, and OpenAI) across 102 physics items, focusing on two main questions: (1) How well do these models perform on conceptual physics tasks involving visual representations? and (2) What are the financial costs associated with their use? The results show high variability in both performance and cost. The performance of the tested models ranges from around 76% to as low as 21%. We also found that expensive models do not always outperform cheaper ones and that, depending on the demands of the context, cheaper models may be sufficiently capable for some tasks. This is especially relevant in contexts where financial resources are limited or for large-scale educational implementation of MLLMs. By providing these analyses, our aim is to inform teachers, institutions, and other educational stakeholders so that they can make evidence-based decisions about the selection of models for use in AI-supported physics education.
@@ 1.1 Large language models in physics education
Large language models (LLMs), which generate human-like text based on statistical patterns in large-scale datasets [6], have significantly accelerated this trend. Their improving performance has sparked growing interest in their application to subject-specific education [7, 8]—including physics, where researchers have assessed various LLMs capabilities using real-world physics and engineering exams from their educational contexts [9, 10, 11, 12].
A growing body of research investigates how LLMs perform across a range of physics education areas. Much of this work has centered on ChatGPT [13], which has demonstrated notable capabilities across multiple physics-related tasks. Studies examining conceptual reasoning [14, 15] and problem solving [16, 17] have shown that the chatbot can produce coherent and well-structured solutions, although it often still exhibits limitations in replicating human-like sensemaking. In essay-style assignments, ChatGPT-generated responses have reached grading levels comparable to high-performing university students [18], raising concerns about the validity of take-home examinations. It has also shown strengths in physics-related programming tasks [10, 19], performing reliably on structured coding tasks, and in lab-based problem solving [20], analyzing experimental data and carrying out statistical analyses.
Very recently, a new class of models—often referred to as reasoning language models (RLMs) [21] (e.g., OpenAI’s o1 and o3, DeepSeek-R1, Alibaba’s QwQ)—has been introduced to specifically enhance performance on complex, multi-step tasks. Unlike earlier chatbots, primarily developed for general linguistic fluency, RLMs are designed to imitate step-by-step reasoning, producing intermediate solution steps, applying domain-relevant procedures, and maintaining better consistency throughout the process [22]. Early evaluations suggest that these models outperform previous generations on benchmark assessments involving mathematical problem solving and STEM-focused conceptual reasoning, leading to a growing interest in their application to physics education [23, 24]. However, these capabilities do not come without important limitations, especially in high-complexity problem solving [25].
@@
In addition, a growing number of research studies in physics education have begun to focus on how AI-based systems can support instructors. One of the reasons is that the increasing student/teacher ratios, largely driven by the persistent shortage of qualified instructors [38], including in physics [39], have placed severe pressure on educational institutions and existing instructors. At the same time, there is growing demand for personalized learning: teachers are expected to deliver timely feedback [40], tailored guidance [41], and fair grading at scale [42], while their capacity is inherently limited [43]. Tasks such as grading [44, 45], tips and feedback generation [46, 47], and support for students with disabilities [48] often involve analysis of student-drawn representations, generation of explanations coupled to such visual representations, and assessing visually complex responses. MLLMs are increasingly considered as one possible solution to offload teachers from these tasks.
@@ 2.1 Model selection
Table 1: Descriptions of the selected AI models. Prices are reported in USD per million tokens. The tested Gemini 2.5 Pro and Flash models are the same as the publicly available and stable versions released in June 2025.
@@
Anthropic [56] Claude Opus 4 claude-opus-4-20250514 15 75
Claude Sonnet 4 claude-sonnet-4-20250514 3 15
Claude Haiku 3.5 claude-3-5-haiku-20241022 0.8 4
Google [57, 58] Gemini 2.5 Pro gemini-2.5-pro-preview-06-05 1.25 10
Gemini 2.5 Flash gemini-2.5-flash-preview-05-20 0.30 2.5
Gemini 2.0 Flash gemini-2.0-flash 0.10 0.4
@@
OpenAI [59] o3 o3-2025-04-16 2 8
o4 mini o4-mini-2025-04-16 1.10 4.4
GPT-4.1 gpt-4.1-2025-04-14 2 8
GPT-4.1 mini gpt-4.1-mini-2025-04-14 0.4 1.6
GPT-4.1 nano gpt-4.1-nano-2025-04-14 0.10 0.4
GPT-4o gpt-4o-2024-11-20 2.5 10
@@
Our selection was guided by the following criteria. First, we included only models that are multimodal and accessible via an application programming interface (API). Second, we aimed to cover a wide performance and pricing spectrum, including both premium-tier systems (e.g., Claude Opus 4, Gemini 2.5 Pro, and o3) and lighter alternatives (e.g., Claude Haiku 3.5, the GPT-4.1 series and Gemini 2.0 Flash). This also allowed us to explore the cost-effectiveness of different models for potential educational deployment.
@@ 2.2 Task selection
To evaluate model performance on image-based conceptual physics tasks, we selected established concept inventories [^1] In order to protect the tests’ integrity, we do not share them. To facilitate the interpretation of this paper, we suggest readers access the tests from the PhysPort website.. Physics concept inventories are research-based multiple-choice assessment tools, developed to probe students’ conceptual understanding of a topic. Each item of the tests typically presents a question and a number of response options. Their standardized structure, focus on conceptual reasoning, often coupled with physics visual representations, makes them well-suited for assessing the capabilities of AI systems in solving visually rich physics tasks.
@@ 2.3 Data collection
Inputs were submitted through official APIs. The temperature parameter was set to 0.7 whenever possible to standardize response variability and reflect a moderate level of generative randomness across conditions [64]. For OpenAI’s reasoning models (o3 and o4-mini), the temperature parameter is not modifiable. A Python-based script was used to automate the process and record responses into json files [65]. Answers (selected letter options appearing at the end of the responses) were extracted into csv files for easier processing. This ensured uniformity across trials and streamlined data collection across models.
@@
This approach was chosen to minimize potential confounding effects introduced by variations in prompt phrasing—a practice that remains largely empirical and difficult to standardize [6]. In particular, we avoided strategies such as few-shot examples [66] or explicit Chain-of-Thought (CoT) [67] cues. As current models often engage in CoT reasoning by default, we focused instead on capturing baseline model behavior.
@@ 2.4 Scoring and analysis
Importantly, this study does not analyze the reasoning or explanations provided in the model outputs. While most responses included extended justifications, our scoring procedure was based solely on the final selected answer. This means we did not assess whether the model’s reasoning was scientifically accurate, coherent, or aligned with its answer. As such, it is entirely possible that a model selected the correct option for incorrect reasons—or vice versa. This design choice reflects the quantitative nature of our analysis, which aims to establish baseline performance levels and relate them to practical considerations such as model cost and accessibility. A full qualitative analysis of reasoning accuracy, coherence, or potential biases would require a different methodological framework and falls outside the scope of this study. For the same reason, we do not offer detailed interpretations of why models performed differently across the four concept inventories. While we propose potential explanations in light of previous research, these remain speculative and are not derived directly from our results.
We determined the cost for running each model by tracking the number of input and output tokes for each model and multiplying them with the listed per-token prices. In reporting the model costs, we normalized the cost to represent the expected cost of running our benchmark on the model a single time—i.e., submitting all 102 items once and receiving one response for each item.
@@ 3.1 Performance analysis
Overall, there is substantial variation in performance across models. o3 and Gemini 2.5 Pro yielded the highest average scores (76.2% and 75.8%, respectively), followed closely by o4 mini (71.5%). Models such as Gemini 2.5 Flash (66.8%), GPT-4.1 mini (53.8%), and GPT-4.1 (52.5%) fall in a middle range, while others like Claude Haiku 3.5 (28.2%), GPT-4.1 nano (25.0%), and Gemma 3-4b (21.0%) scored considerably lower. The range between the highest and lowest average model score exceeds 50 percentage points, highlighting the breadth of performance observed across the models.
@@
Avg AI (per concept inventory) Perf 64.5 62.9 41.8 35.2
SD 4.6 5.5 5.8 6.2
SEM 1.4 1.7 1.8 1.9
@@
Performance also varied across concept inventories (Fig. 1). The average performance across all tested models on BEMA (64.5%) and TUG-K (62.9%) was higher than on QMVI (41.8%) and FTGOT (35.2%). These patterns are consistent across most models. For example, o3 and Gemini 2.5 Pro exceeded 84% on both BEMA and TUG-K, while none of them surpassed 75% on QMVI or 52% on FTGOT. Across inventories, the maximum difference between a given model’s highest and lowest test score ranged from 30 to 50 percentage points, suggesting some inventories (especially the FTGOT) have higher difficulty even for high-performing models.
Figure 1: Distribution of model scores across the selected inventories, together with the average distribution (right-most violin). Individual models are color-coded according to their developer. The full set of data is available in Table 3. For reference, post-instruction student performance is marked with a white diamond. Student data from [68, 69, 70, 71].
The statistical measures provide additional detail. SDs were mostly between 3 and 7 percentage points of the total available points, with most SEM values below 2.5%. In many cases, higher-scoring models (e.g., GPT-4o, Gemini 2.5 Pro) exhibited both higher average performance and lower spread of scores, indicating greater consistency across runs. Some lower-performing models, such as GPT-4.1 nano and Gemma 3-4b, also showed relatively low variability of scores, suggesting consistent but inaccurate responses. In contrast, several mid-range models displayed greater variability, reflecting a mix of correct and incorrect answers across trials.
Looking at performance within model families, some patterns emerge. Among the OpenAI models, o3 and o4 mini performed better than GPT-4.1 and GPT-4.1 mini, while GPT-4.1 nano scored lowest. Similarly, for Google’s models, Gemini 2.5 Pro outperformed Gemini 2.5 Flash, and Gemini 2.5 Flash outperformed the earlier Gemini 2.0 Flash, as well as Gemma. For Anthropic, Claude Opus 4 performed better than Claude Sonnet 4, while Claude Haiku 3.5 scored lowest in that family. While exact differences vary, in most cases each successive tier within a provider correlates with a performance increase of 10–20 percentage points.
@@ 3.2 Cost analysis
To better understand the relationship between model performance and usage cost, we plotted each model’s average score against its cost. The resulting plot in Fig. 2 reveals a few patterns.
@@
First, while there is an overall correlation between cost and performance, the relationship is not proportional, and there are some interesting outliers. Some of the top-performing models, such as o3 and o4 mini, are positioned in the cost range between \$1.00 and \$1.50. Despite not being the most expensive, they achieved accuracy scores of 76.2% and 71.5%, respectively, which is comparable to the much more costly (\$4.68) Gemini 2.5 Pro model, with a performance of 75.8%. On the other hand, Claude Opus 4, only marginally cheaper than Gemini 2.5 Pro, achieved only 57.0% on the benchmark, significantly underperforming relative to its cost.
@@
Third, several models strike a compelling balance between affordability and capability. Gemini 2.5 Flash, when the “reasoning is enabled”, performs in the high 60% range while maintaining the cost just below \$1.00. Even more interesting is the performance of Gemini 2.5 Flash with “reasoning disabled”. It achieved a performance in the lower 60% range at the cost of only \$0.31, making it an interesting contender for low-cost uses. For less demanding uses, GPT 4.1 mini might be a good candidate. Interestingly, it outperformed its bigger brother, GPT-4.1, and displayed performance in the mid-50% range at the price of only \$0.10. Note that this is 45-times cheaper than Claude Opus 4, with essentially the same level of performance.
Finally, grouping by provider reveals broader trends. OpenAI’s models tend to cluster in the upper-left quadrant—combining solid performance with reasonable costs. Google’s lineup is more varied, with models ranging from the very strong but expensive Gemini 2.5 Pro, to the underperforming but completely free Gemma models. Anthropic’s Claude models fall in the middle or lower performance tiers, despite being relatively expensive, suggesting a less favorable price-to-performance ratio for this task set.
@@
$Token_{in}$ $Cost_{in}$ $Token_{reas}$ $Cost_{reas}$ $Token_{out}$ $Cost_{out}$
o3 88152 0.1763 159769.6 1.2782 165696.1 1.3256 1.5019
Gemini 2.5 Pro 30600 0.0383 403024.3 4.0302 61533.2 0.6153 4.6838
o4 mini 211157 0.2323 168288 0.7405 176051.2 0.7746 1.0069
Gemini 2.5 Flash 30600 0.0092 264592.4 0.6615 71755.5 0.1794 0.8500
Gemini 2.5 Flash (no reasoning) 30600 0.0092 122266.9 0.3057 0.3148
Claude Opus 4 138554 2.0783 32811.6 2.4609 4.5392
GPT-4.1 mini 159056 0.0636 19758.5 0.0316 0.0952
Claude Sonnet 4 138554 0.4157 37799.4 0.5670 0.9827
GPT-4.1 99314 0.1986 14777.1 0.1182 0.3168
Gemini 2.0 Flash 231738 0.0232 9836.9 0.0040 0.0271
GPT-4o 99314 0.2483 15584.6 0.1558 0.4041
@@
Claude Haiku 3.5 138702 0.1110 13323.6 0.0533 0.1643
GPT-4.1 nano 299810 0.0300 12009.3 0.0048 0.0348
@@
Overall, these results suggest that model selection for educational applications should be guided by empirical performance data rather than assumptions based on provider reputation or listed per-token prices. Lower-cost models may in some cases offer equal, comparable, or even superior performance on physics conceptual tasks involving visual representations.

### v2

@@ Abstract
Multimodal large language models (MLLMs) capable of processing both text and visual inputs are increasingly being explored for uses in physics education, such as tutoring, formative assessment, and grading. This study evaluates a range of publicly available MLLMs on a set of standardized, image-based physics research-based conceptual assessments (concept inventories). We benchmark 15 models from three major providers (Anthropic, Google, and OpenAI) across 102 physics items, focusing on two main questions: (1) How well do these models perform on conceptual physics tasks involving visual representations? and (2) What are the financial costs associated with their use? The results show high variability in both performance and cost. The performance of the tested models ranges from 81.5% to as low as 21%. We also found that expensive models do not always outperform cheaper ones and that, depending on the demands of the context, cheaper models may be sufficiently capable for some tasks. This is especially relevant in contexts where financial resources are limited or for large-scale educational implementation of MLLMs. By providing these analyses, our aim is to inform teachers, institutions, and other educational stakeholders so that they can make evidence-based decisions about the selection of models for use in AI-supported physics education.
@@ 1.1 Large language models in physics education
Large language models (LLMs), which generate human-like text based on statistical patterns in large-scale datasets [6], have significantly accelerated this trend. Their improving performance has sparked growing interest in their application to subject-specific education [7, 8]—including physics, where researchers have assessed various LLMs capabilities using real-world physics and engineering exams from their educational contexts [9, 10, 11, 12, 13].
A growing body of research investigates how LLMs perform across a range of physics education areas. Much of this work has centered on ChatGPT [14], which has demonstrated notable capabilities across multiple physics-related tasks. Studies examining conceptual reasoning [15, 16] and problem solving [17, 18, 19] have shown that the chatbot can produce coherent and well-structured solutions, although it often still exhibits limitations in replicating human-like sensemaking. In essay-style assignments, ChatGPT-generated responses have reached grading levels comparable to high-performing university students [20], raising concerns about the validity of take-home examinations. It has also shown strengths in physics-related programming tasks [12, 21], performing reliably on structured coding tasks, and in lab-based problem solving [22, 23], analyzing experimental data and carrying out statistical analyses.
Very recently, a new class of models—often referred to as reasoning language models (RLMs) [24] (e.g., OpenAI’s o3 and GPT-5, DeepSeek-R1, Alibaba’s QwQ)—has been introduced to specifically enhance performance on complex, multi-step tasks. Unlike earlier chatbots, primarily developed for general linguistic fluency, RLMs are designed to imitate step-by-step reasoning, producing intermediate solution steps, applying domain-relevant procedures, and maintaining better consistency throughout the process [25]. Early evaluations suggest that these models outperform previous generations on benchmark assessments involving mathematical problem solving and STEM-focused conceptual reasoning, leading to a growing interest in their application to physics education [26, 27]. However, these capabilities do not come without important limitations, especially in high-complexity problem solving [28].
@@
In addition, a growing number of research studies in physics education have begun to focus on how AI-based systems can support instructors. One of the reasons is that the increasing student/teacher ratios, largely driven by the persistent shortage of qualified instructors [41], including in physics [42], have placed severe pressure on educational institutions and existing instructors. At the same time, there is growing demand for personalized learning: teachers are expected to deliver timely feedback [43], tailored guidance [44], and fair grading at scale [45], while their capacity is inherently limited [46]. Tasks such as grading [47, 48, 49, 50], tips and feedback generation [51, 52, 53], and support for students with disabilities [54] often involve analysis of student-drawn representations, generation of explanations coupled to such visual representations, and assessing visually complex responses. MLLMs are increasingly considered as one possible solution to offload teachers from these tasks.
@@ 2.1 Model selection
Table 1: Descriptions of the selected AI models. Prices are reported in USD per million tokens.
@@
Anthropic [62] Claude Opus 4 claude-opus-4-20250514 15.00 75.00
Claude Sonnet 4 claude-sonnet-4-20250514 3.00 15.00
Claude Haiku 3.5 claude-3-5-haiku-20241022 0.80 4.00
Google [63, 64] Gemini 2.5 Pro gemini-2.5-pro-preview-06-05 1.25 10.00
Gemini 2.5 Flash gemini-2.5-flash-preview-05-20 0.30 2.50
Gemini 2.0 Flash gemini-2.0-flash 0.10 0.40
@@
OpenAI [65] GPT-5 gpt-5-2025-08-07 1.25 10.00
GPT-5 mini gpt-5-mini-2025-08-07 0.25 2.00
GPT-5 nano gpt-5-nano-2025-08-07 0.05 0.40
o3 o3-2025-04-16 2.00 8.00
o4 mini o4-mini-2025-04-16 1.10 4.40
GPT-4.1 gpt-4.1-2025-04-14 2.00 8.00
GPT-4.1 mini gpt-4.1-mini-2025-04-14 0.40 1.60
GPT-4.1 nano gpt-4.1-nano-2025-04-14 0.10 0.40
GPT-4o gpt-4o-2024-11-20 2.50 10.00
@@
Our selection was guided by the following criteria. First, we included only models that are multimodal and accessible via an application programming interface (API). Second, we aimed to cover a wide performance and pricing spectrum, including both premium-tier systems (e.g., Claude Opus 4, Gemini 2.5 Pro, GPT-5 and o3) and lighter alternatives (e.g., Claude Haiku 3.5, GPT-5 mini, GPT-5 nano, the GPT-4.1 series, and Gemini 2.0 Flash). This also allowed us to explore the cost-effectiveness of different models for potential educational deployment.
@@ 2.2 Task selection
To evaluate model performance on image-based conceptual physics tasks, we selected established concept inventories [^2] In order to protect the tests’ integrity, we do not share them. To facilitate the interpretation of this paper, we suggest readers access the tests from the PhysPort website.. Physics concept inventories are research-based multiple-choice assessment tools, developed to probe students’ conceptual understanding of a topic. Each item of the tests typically presents a question and a number of response options. Their standardized structure, focus on conceptual reasoning, often coupled with physics visual representations, makes them well-suited for assessing the capabilities of AI systems in solving visually rich physics tasks. Fig. 1 shows an example of a test item as it was submitted.
Figure 1: An example of one of the 102 test item screenshots submitted to the models. It consists of textual and image parts, as well as a multiple-choice answer list. Image adapted from [33] under the $C$-BY 4.0 license.
@@ 2.3 Data collection
Inputs were submitted through official APIs. The temperature parameter was set to 0.7 whenever possible to standardize response variability and reflect a moderate level of generative randomness across conditions [70]. For OpenAI’s reasoning models (o3 and o4-mini), the temperature parameter is not modifiable. For GPT-5 series models, the parameter “reasoning effort” was set to medium. A Python-based script was used to automate the process and record responses into json files [71]. Answers (selected letter options appearing at the end of the responses) were extracted into csv files for easier processing. This ensured uniformity across trials and streamlined data collection across models.
@@
This approach was chosen to minimize potential confounding effects introduced by variations in prompt phrasing—a practice that remains largely empirical and difficult to standardize [6]. In particular, we avoided strategies such as few-shot examples [72] or explicit Chain-of-Thought (CoT) cues [73]. As current models often engage in CoT reasoning by default, we focused instead on capturing baseline model behavior.
@@ 2.4 Scoring and analysis
Importantly, this study does not analyze the reasoning or explanations provided in the model outputs. While responses often included extended justifications, our scoring procedure was based solely on the final selected answer. This means we did not assess whether the model’s reasoning was scientifically accurate, coherent, or aligned with its answer. As such, it is entirely possible that a model selected the correct option for incorrect reasons—or vice versa. This design choice reflects the quantitative nature of our analysis, which aims to establish baseline performance levels and relate them to practical considerations such as model cost and accessibility. A full qualitative analysis of reasoning accuracy, coherence, or potential biases would require a different methodological framework and falls outside the scope of this study. For the same reason, we do not offer detailed interpretations of why models performed differently across the four concept inventories. While we propose potential explanations in light of previous research, these remain speculative and are not derived directly from our results.
We determined the cost for running each model by tracking the number of input and output tokens for each model and multiplying them by the listed per-token prices. In reporting the model costs, we normalized the cost to represent the expected cost of running our benchmark on the model a single time—i.e., submitting all 102 items once and receiving one response for each item. The costs reported in this paper include both input and output costs. Calculating the expected input costs was straightforward; we simply multiplied the number of tokens for each input (which was the same for all 10 iterations of an item on a given model) with the model-specific input price per token (see Table 1 for token-based prices). For calculating the output costs, we took the average number of tokens outputted by a model for a given item across all 10 iterations, and summed these item averages across all items for each model to obtain the expected number of tokens for the entire benchmark. Multiplying this number with the price per output token for the model gave us the expected output cost of running the entire benchmark for that model. For reasoning models, reasoning tokens are counted and charged as output tokens, even though they are not directly accessible to the user. In these models, reasoning tokens represented most of the cost.
@@ 3.1 Performance analysis
Overall, there is substantial variation in performance across models. GPT-5 was the best performing model (81.5%). o3, Gemini 2.5 Pro and GPT-5 mini yielded high average scores (76.2%, 75.8%, and 75.0% respectively), followed closely by o4 mini (71.5%). Models such as Gemini 2.5 Flash (66.8%), GPT-5 nano (60.7%), Claude Opus 4 (57.0%), GPT-4.1 mini (53.8%), Claude Sonnet 4 (53.7%), and GPT-4.1 (52.5%) fall in the middle range. Others, like Claude Haiku 3.5 (28.2%), GPT-4.1 nano (25.0%), and Gemma 3-4b (21.0%) scored considerably lower. The range between the highest and lowest average model score exceeds 60 percentage points, highlighting the breadth of performance observed across the models.
@@
Avg AI (per concept inventory) Perf 66.7 64.4 44.4 35.1
SD 4.6 5.5 5.7 6.3
SEM 1.4 1.7 1.8 2.0
@@
Performance also varied across concept inventories (Fig. 2). The average performance across all tested models on BEMA (66.7%) and TUG-K (64.4%) was higher than on QMVI (44.4%) and FTGOT (35.1%). These patterns are consistent across most models. For example, GPT-5 performed above 90% for BEMA and TUG-K, about 10 percentage points lower on QMVI, and less than 50% on FTGOT. Following a similar pattern, o3 and Gemini 2.5 Pro exceeded 84% on both BEMA and TUG-K, while none of them surpassed 75% on QMVI or 52% on FTGOT. Across inventories, the maximum difference between a given model’s highest and lowest test score ranged from 30 to 50 percentage points, suggesting some inventories (especially FTGOT) have higher difficulty even for high-performing models.
Figure 2: Distribution of model scores across the selected inventories, together with the average distribution (right-most violin). Individual models are color-coded according to their developer. The full set of data is available in Table 3. For reference, post-instruction university student performance is marked with a white diamond. Student data from [74, 75, 68, 67].
The statistical measures provide additional detail. SDs were mostly between 3 and 7 percentage points of the total available points, with most SEM values below 2.5%. In many cases, higher-scoring models exhibited both higher average performance and lower spread of scores, indicating greater consistency across runs. Some lower-performing models also showed relatively low variability of scores, suggesting consistent but inaccurate responses. In contrast, several mid-range models displayed greater variability, reflecting a mix of correct and incorrect answers across trials.
Looking at performance within model families, some patterns emerge. Among the OpenAI models, GPT-5, o3, GPT-5 mini, o4 mini and even GPT-5 nano, performed better than GPT-4.1 and GPT-4.1 mini, while GPT-4.1 nano scored lowest. Similarly, for Google’s models, Gemini 2.5 Pro outperformed Gemini 2.5 Flash, and Gemini 2.5 Flash outperformed the earlier Gemini 2.0 Flash, as well as Gemma. For Anthropic, Claude Opus 4 performed better than Claude Sonnet 4, while Claude Haiku 3.5 scored lowest in that family. While exact differences vary, in most cases each successive tier within a provider correlates with a performance increase of 10–20 percentage points.
@@ 3.2 Cost analysis
To better understand the relationship between model performance and usage cost, we plotted each model’s average score against its cost. The resulting plot in Fig. 3 reveals a few patterns. For a more detailed breakdown of the costs, see Table 4.
@@
First, while there is an overall correlation between cost and performance, the relationship is not proportional, and there are some interesting outliers. GPT-5 is the best performing model with an accuracy score of 81.5% at the cost of \$2.18. Some of the other relatively well-performing models, such as o3 and o4 mini, are positioned in the cost range between \$1.00 and \$1.50. Despite not being the most expensive, they achieved accuracy scores of 76.2% and 71.5%, respectively, which is comparable to the much more costly Gemini 2.5 Pro model (\$4.68), with a performance of 75.8%. The clear outlier here is GPT-5 mini, with the accuracy score of 75% at an outstandingly low cost of \$0.26. On the other hand, Claude Opus 4, only marginally cheaper than Gemini 2.5 Pro, achieved only 57.0% on the benchmark, significantly underperforming relative to its cost.
@@
Third, several models strike a compelling balance between affordability and capability. Gemini 2.5 Flash, when the “reasoning is enabled”, performs in the high 60% range while maintaining the cost just below \$1.00. Even more interesting is the performance of Gemini 2.5 Flash with “reasoning disabled”. It achieved a performance in the lower 60% range at the cost of only \$0.31, making it an interesting contender for low-cost uses. For less demanding uses or tight budgets, GPT-5 nano appears to be a good candidate, with low cost and performance just above 60%. Note that this is 38 times cheaper than Claude Opus 4, which performs at only 57%. Interestingly, GPT-4.1 mini outperformed its bigger brother, GPT-4.1, and displayed performance in the mid-50% range at the price of only \$0.10. Once again, it is notable that GPT-5 mini is in a class of its own at this price point, performing about 10 percentage points better than other models at the cost of \$0.27.
Finally, grouping by provider reveals broader trends. OpenAI’s models tend to cluster in the upper-left quadrant—combining solid performance with reasonable costs. Google’s lineup is more varied, with models ranging from the very strong but expensive Gemini 2.5 Pro, to the underperforming but completely free Gemma models. Anthropic’s Claude models fall in the middle or lower performance tiers, despite being relatively expensive, suggesting a less favorable price-to-performance ratio for this particular task set.
@@
$Token_{in}$ $Cost_{in}$ $Token_{rea}$ $Cost_{rea}$ $Token_{out}$ $Cost_{out}$
GPT-5 82622 0.103 204237 2.042 4016 0.040 2.186
o3 88152 0.176 159770 1.278 5927 0.047 1.502
Gemini 2.5 Pro 30600 0.038 403024 4.030 61533 0.615 4.684
GPT-5 mini 148923 0.037 112128 0.224 3949 0.008 0.269
o4 mini 211157 0.232 168288 0.741 7763 0.034 1.007
Gemini 2.5 Flash 30600 0.009 264592 0.662 71756 0.179 0.850
Gemini 2.5 Flash (no reasoning) 30600 0.009 122267 0.306 0.315
GPT-5 nano 184807 0.009 278515 0.111 2880 0.001 0.122
Claude Opus 4 138554 2.078 32812 2.461 4.539
GPT-4.1 mini 159056 0.064 19759 0.032 0.095
Claude Sonnet 4 138554 0.416 37799 0.567 0.983
GPT-4.1 99314 0.199 14777 0.118 0.317
Gemini 2.0 Flash 231738 0.023 9837 0.004 0.027
GPT-4o 99314 0.248 15585 0.156 0.404
@@
Claude Haiku 3.5 138702 0.111 13324 0.053 0.164
GPT-4.1 nano 299810 0.030 12009 0.005 0.035
@@
Overall, these results suggest that model selection for educational applications should be guided by empirical performance data rather than assumptions based on provider reputation or listed per-token prices. Lower-cost models may, in some cases, offer equal, comparable, or even superior performance on physics conceptual tasks involving visual representations.